# Efficient Design for Multi-user Downlink Beamforming with Reconfigurable Intelligent Surface

### v2

@@ V Generalization to RIS-Assisted Multi-user Downlink Beamforming
The weights $\{a_{i}\}$ for all $G$ users can be derived in closed-form. This is because when the transmit beamformer can be individually designed for each user, at the optimality, all SINR constraints in (5b) is attained with equality. Specifically, let ${\bf c}_{i}\triangleq{\bf R}^{-1}(\boldsymbol{\lambda})\widetilde{{\bf h}}_{i1}$, $i\in{\cal G}$, and we have the following the SINR equality constraint
$\frac{a_{i}^{2}}{\gamma_{i1}}|{\bf c}_{i}^{H}{\bf h}_{i1}|^{2}=\!\!\sum_{j\in{\cal G},j\neq i}\!\!a_{j}^{2}|{\bf c}_{j}^{H}{\bf h}_{j1}|^{2}+\sigma^{2},\quad\text{for}~{}i\in{\cal G}.$ (22)
Thus, $\{a_{i}^{2}\}$ can be obtained by solving the above $G$ linear equations. Define the $G\times G$ matrix ${\bf V}$ with its $ij$-th entry as
$[{\bf V}]_{ij}=\begin{cases}\frac{1}{\gamma_{i1}}|{\bf c}_{i}^{H}{\bf h}_{i1}|^{2},&j=i\\ -|{\bf c}_{i}^{H}{\bf h}_{j1}|^{2},&j\neq i.\end{cases}$ (23)
Define $\tilde{{\bf a}}=[a_{1}^{2},...,a_{G}^{2}]^{T}$. Then, the solution to the $G$ linear equations in (22) can be compactly written as
$\tilde{{\bf a}}=\sigma^{2}{\bf V}^{-1}\mathbf{1}$ (24)
where $\mathbf{1}$ is an all-one vector.
Finally, without the loss of optimality, we obtain
$a_{i}=\sqrt{[\tilde{{\bf a}}]_{i}},\quad i\in{\cal G}.$ (25)
Thus, the joint optimization of $({\bf w},{\bf e})$ for ${\cal P}_{o}$ by the AMBF approach involves alternating between computing the closed-form solution ${\bf w}_{i}$ in (21) and solving $\widetilde{{\cal S}}_{{\bf w}}$ for ${\bf e}$ using the updates in (17) and (18).
Remark 3.
RIS-assisted downlink multi-user beamforming design for the unicast scenario has been studied in many recent works. However, only a few of them considered the QoS problem for transmit power minimization [5, 21]. Perhaps partially this is due to the difficulty in using AO effectively to solve the joint BS and RIS beamforming problem. In [5], directly applying AO to the joint BS and RIS beamforming to minimize the BS transmit power is considered. As discussed at the beginning of Section III, direct AO renders the RIS beamforming a feasibility problem, which can not provide an effective solution. Thus, some modification to the feasibility problem is proposed to find a better solution, and a low-complexity heuristic method based on the sum of channel gains is also proposed. The joint beamforming design for RIS-assisted non-orthogonal multiple access (NOMA) networks is considered in [21], where the direct AO approach is also adopted, with the RIS beamforming being treated as a feasibility problem. Both [5] and [21] adopt SDP or SDR to solve the corresponding alternating subproblems. In our work, we first transform the QoS problem into ${\cal P}_{2}$ in (III-A). It allows us to break the problem into two subproblems, for which AO is effective in obtaining a good solution. The subproblems can be solved by the closed-form solution in (21) and the fast algorithm in Section IV-B, which is a scalable solution for large-scale systems.
VI RIS-Assisted Multicast MMF Problem
We now consider the weighted max-min fair (MMF) design for the RIS-assisted multicast scenario. The objective is to maximize the minimum weighted SINR among users, subject to the BS transmit power budget and the unit-modulus constraint for the RIS elements. The weighted SINR is the ratio between SINR and a pre-specified SINR target, which sets certain fairness among users. This joint optimization problem w.r.t. ${\bf w}$ and ${\bf e}$ is formulated as
@@ VI RIS-Assisted Multicast MMF Problem
where $P$ is the transmit power budget. Similar to the QoS problem ${\cal P}_{o}$, problem ${\cal S}_{o}$ is also non-convex and NP-hard.
For the conventional downlink multicast beamforming design at the BS, it is known that the QoS and MMF problems are inverse problems [28]: the role of the objective function and the constraints are switched in the two problems, and we can solve the MMF problem by iteratively solving the QoS problem along with a bi-section search over the minimum SINR target for the QoS problem until the total transmit power meets the power budget in the MMF problem. Naturally, we want to know whether such a relation also holds for the RIS-assisted multicast beamforming design, which now also includes the RIS passive beamformer optimization, in addition to the BS transmit beamformers. We show below that this is indeed the case.
@@
For RIS-assisted multicast beamforming, the QoS problem ${\cal P}_{o}$ and the MMF problem ${\cal S}_{o}$ are inverse problems. In particular, we explicitly parameterize ${\cal S}_{o}$ as ${\cal S}_{o}(\boldsymbol{\gamma},P)$ for given $\boldsymbol{\gamma}$ and $P$ with the maximum objective value denoted as $t^{o}={\cal S}_{o}(\boldsymbol{\gamma},P)$, and we also parameterize ${\cal P}_{o}$ as ${\cal P}_{o}(\boldsymbol{\gamma})$ with $\boldsymbol{\gamma}$ being the SINR target vector, and the minimum power obtained as $P={\cal P}_{o}(\boldsymbol{\gamma})$. Then, ${\cal P}_{o}(\boldsymbol{\gamma})$ and ${\cal S}_{o}(\boldsymbol{\gamma},P)$ have the following relations:
@@
Based on the above relation, we can solve the MMF problem ${\cal S}_{o}$ using Algorithm 1. In particular, by the relation in (27), ${\cal S}_{o}$ can be solved by solving its inverse QoS problem ${\cal P}_{o}$ iteratively along with a bi-section search for $t^{o}$. However, since the size of the RIS reflection elements is typically large $M\gg 1$, iteratively solving ${\cal P}_{o}$ will incur relatively high computational complexity. We need a more computationally efficient algorithm to solve the large-scale joint optimization problem ${\cal S}_{o}$. Aiming at this goal, we present a fast algorithm to solve ${\cal S}_{o}$ directly.
VI-A Proposed Fast Algorithm for Solving ${\cal S}_{o}$
For the conventional downlink scenario without RIS, as mentioned in Section IV-B, PSA is proposed in [12] to obtain a near-stationary solution for the multicast beamforming MMF problem. Below, we show that we can adopt PSA to solve the MMF problem efficiently for the RIS-assisted multicast scenario directly as well.
From Proposition 1, problems ${\cal P}_{o}$ and ${\cal S}_{o}$ are the inverse problems. This means the optimal structure of the BS multicast beamformer ${\bf w}$ still has the form given in (13). Indeed, with given ${\bf e}$, if we treat the RIS-assisted channel between the BS and user $k$ in group $i$ as the equivalent channel $\tilde{{\bf h}}_{ik}$, ${\cal S}_{o}$ reduces to the conventional multicast MMF problem, for which the optimal beamforming structure is (13) [11]. However, the challenge for the MMF problem is that computing $\boldsymbol{\lambda}$ in ${\bf R}(\boldsymbol{\lambda})$ is not straightforward. Fortunately, the asymptotic expression of ${\bf R}(\boldsymbol{\lambda})$ for large $N$ is obtained in closed form in [11], which can be used as an approximate expression of ${\bf R}(\boldsymbol{\lambda})$ to further simplify the computation. Specifically, let $\tilde{{\bf h}}_{iK}=\sqrt{\tilde{\beta}_{ik}}{\bf g}_{ik}$ where $\tilde{\beta}_{ik}$ represents the large-scale channel variance and $\tilde{{\bf g}}_{ik}\sim{\cal C}{\cal N}(\mathbf{0},{\bf I})$. Then, the asymptotic expression of ${\bf R}(\boldsymbol{\lambda})$ has the following simple closed form
@@ VI-A Proposed Fast Algorithm for Solving ${\cal S}_{o}$
${\bf R}(\boldsymbol{\lambda})\approx{\bf I}+\frac{P\bar{\beta}_{{\bf h}}}{\sigma^{2}K_{\text{tot}}}\sum_{i=1}^{G}\sum_{k=1}^{K_{i}}{\bf g}_{ik}{\bf g}_{ik}^{H}\triangleq\widetilde{{\bf R}}$ (28)
@@
where $\bar{\beta}_{{\bf h}}\triangleq 1/\Big{(}\frac{1}{K_{\text{tot}}}\sum_{i=1}^{G}\sum_{k=1}^{K_{i}}\frac{1}{\tilde{\beta}_{ik}}\Big{)}$ is the harmonic mean of the large-scale channel variance of all users. Using $\widetilde{{\bf R}}$ in (28), we can replace ${\bf w}_{i}$ in (13) with the following expression:
${\bf w}_{i}\,\approx\,\widetilde{{\bf R}}^{-1}\widetilde{{\bf H}}_{i}{\bf a}_{i}\,,\;i\in{\cal G}.$ (29)
Below, we develop our fast algorithm to solve ${\cal S}_{o}$ based on ${\bf w}_{i}$ in (29) to reduce the computational complexity.
@@
Using ${\bf w}_{i}$ in (29), we first convert ${\cal S}_{o}$ into a joint optimization problem w.r.t. $\{{\bf a}_{i}\}$ and ${\bf e}$, given by
@@
where ${\bf a}\triangleq[{\bf a}_{1}^{H},\cdots,{\bf a}_{G}^{H}]^{H}$, $\widetilde{{\bf C}}_{i}\triangleq\widetilde{{\bf R}}^{-1}\widetilde{{\bf H}}_{i}$, and the constraints in (26c) are replaced by the equivalent set of constraints in (30c) and (30d) based on (9). Similar to $\widetilde{{\cal S}}_{{\bf w}}$ in (III-B2), we relax $S_{o}^{\text{eq}}$ into the following problem by moving the equality constraint in (30d) into the objective function as a penalty term with a penalty weight $\delta>0$:
$\widetilde{{\cal S}}_{o}:\max_{{\bf a},{\bf e}}$ $\ \bigg{(}\!\min_{i,k}\ \ \frac{\textrm{SINR}_{ik}}{\gamma_{ik}}\bigg{)}+\delta\frac{{\bf e}^{H}{\bf e}}{M}$ (31a)
@@
Problem $\widetilde{{\cal S}}_{o}$ is similar to $\widetilde{{\cal S}}_{\bf w}$ in (III-B2), except that $\widetilde{{\cal S}}_{o}$ is a joint optimization problem for $({\bf a},{\bf e})$ while $\widetilde{{\cal S}}_{\bf w}$ is for ${\bf e}$ only. Nonetheless, we can still apply PSA to compute a near-stationary joint solution $({\bf a},{\bf e})$ for $\widetilde{{\cal S}}_{o}$ efficiently, similar to Section IV-B, as shown below.
VI-A1 Problem Reformulation
@@
Let ${\bf x}\triangleq[{\bf a}^{H},{\bf e}^{H}]^{H}$. By substituting the SINR expression in (2) into the objective function in (31a), we define
@@
$\varphi_{ik}({\bf x})\triangleq$ $-\!\frac{1}{\gamma_{ik}}\frac{|{\bf a}_{i}^{H}\widetilde{{\bf C}}_{i}^{H}\!({\bf h}_{ik}^{\text{d}}\!\!+\!{\bf G}_{ik}{\bf e})|^{2}}{\sum_{j\neq i}|{\bf a}_{j}^{H}\widetilde{{\bf C}}_{j}^{H}\!({\bf h}_{ik}^{\text{d}}\!\!+\!{\bf G}_{ik}{\bf e})|^{2}\!+\!\sigma^{2}}-\delta\frac{{\bf e}^{H}{\bf e}}{M},$ (32)
@@
for $k\in{\cal K}_{i}$, $i\in{\cal G}$. Following the similar arguments from (14)–(16), let $\boldsymbol{\varphi}({\bf x})$ be a $K_{\text{tot}}\times 1$ vector containing all $\varphi_{ik}({\bf x})$’s, and let ${\bf y}$ denote a $K_{\text{tot}}\times 1$ vector with ${\bf y}\succcurlyeq 0$ and ${\bf 1}^{T}{\bf y}=1$. Then, $\widetilde{{\cal S}}_{o}$ can be transform into the following equivalent problem
$\widetilde{{\cal S}}_{o}^{\prime}:\min_{{\bf x}\in{\cal U}}\max_{{\bf y}\in{\cal Y}}$ $\ g({\bf x},{\bf y})$ (33)
@@
where $g({\bf x},{\bf y})\triangleq\boldsymbol{\varphi}^{T}({\bf x}){\bf y}$, ${\cal Y}\triangleq\{{\bf y}:{\bf y}\succcurlyeq 0,\mathbf{1}^{T}{\bf y}=1\}$, and ${\cal U}\triangleq\{{\bf x}:|e_{m}|^{2}\leq 1,\ m\in{\cal M};\ \sum_{i=1}^{G}\|\widetilde{{\bf C}}_{i}{\bf a}_{i}\|^{2}\leq P\}$. Both ${\cal Y}$ and ${\cal U}$ are compact convex sets.
VI-A2 The Projected Subgradient Algorithm

### v3

@@ IV-D Special Case $K_{i}=1$
The weights $a_{i}$’s for all $G$ users can be derived in closed-form in this unicast case. This is because when the transmit beamformer can be individually designed for each user, at the optimality, all SINR constraints in (4b) can be attained with equality. Denote ${\bf c}_{i}\triangleq{\bf R}^{-1}(\boldsymbol{\lambda})\widetilde{{\bf h}}_{i}$, $i\in{\cal G}$. Then, we have the following SINR equality constraints
$\frac{a_{i}^{2}}{\gamma_{i}}|{\bf c}_{i}^{H}{\bf h}_{i}|^{2}=\!\!\sum_{j\in{\cal G},j\neq i}\!\!a_{j}^{2}|{\bf c}_{j}^{H}{\bf h}_{i}|^{2}+\sigma^{2},\quad\text{for}~{}i\in{\cal G}.$ (23)
Solving the above $G$ linear equations, we obtain $a_{i}^{2}$, for $i\in{\cal G}$. In a matrix representation, define a $G\times G$ matrix ${\bf V}$ with its $ij$-th entry as
$[{\bf V}]_{ij}=\begin{cases}\frac{1}{\gamma_{i}}|{\bf c}_{i}^{H}{\bf h}_{i}|^{2},&j=i\\ -|{\bf c}_{j}^{H}{\bf h}_{i}|^{2},&j\neq i.\end{cases}$ (24)
Then, the solution to the $G$ linear equations in (23) can be compactly written as
$\tilde{{\bf a}}^{\text{\tiny sq}}=\sigma^{2}{\bf V}^{-1}\mathbf{1}$ (25)
where $\tilde{{\bf a}}^{\text{\tiny sq}}=[a_{1}^{2},...,a_{G}^{2}]^{T}$, and $\mathbf{1}$ is an all-one vector. Finally, without the loss of optimality, we have $a_{i}=\sqrt{[\tilde{{\bf a}}]_{i}}$, $i\in{\cal G}$.
In summary, in this special unicast case, the AMBF approach for solving the joint optimization of $({\bf w},{\bf e})$ for ${\cal P}_{o}$ involves alternating between computing the closed-form solution ${\bf w}_{i}$ in (22) and solving $\widetilde{{\cal S}}_{{\bf w}}$ for ${\bf e}$ using the updates in (18) and (19), which is a computationally efficient solution.
V RIS-Assisted Multicast MMF Problem
Another commonly considered multicast beamforming formulation is the weighted MMF problem. The objective is to maximize the minimum weighted SINR among users, subject to the BS transmit power budget and the unit-modulus constraint for the RIS elements. This joint optimization problem w.r.t. ${\bf w}$ and ${\bf e}$ is formulated as
@@ V RIS-Assisted Multicast MMF Problem
where $P$ is the maximum transmit power limit, and the weight $\gamma_{ik}$ is a pre-specified SINR target, which can be used to set certain fairness among users.
Similar to the QoS problem ${\cal P}_{o}$, the MMF problem ${\cal S}_{o}$ is non-convex NP-hard. We are interested in the relationship between ${\cal P}_{o}$ and ${\cal S}_{o}$. Note that the RIS-assisted multicast beamforming design includes optimization of the RIS passive beamformer in addition to the BS transmit beamformers. We show below that these two problems are inverse problems.
@@
For RIS-assisted multicast beamforming, the QoS problem ${\cal P}_{o}$ and the MMF problem ${\cal S}_{o}$ are inverse problems. Specifically, let $\boldsymbol{\gamma}$ be the vector containing all $\gamma_{ik}$’s. Parameterize ${\cal S}_{o}$ as ${\cal S}_{o}(\boldsymbol{\gamma},P)$ for given $\boldsymbol{\gamma}$ and $P$ with the maximum objective value denoted as $t^{o}={\cal S}_{o}(\boldsymbol{\gamma},P)$, and parameterize ${\cal P}_{o}$ as ${\cal P}_{o}(\boldsymbol{\gamma})$ with $\boldsymbol{\gamma}$ being the SINR target vector, with the minimum power objective value as $P={\cal P}_{o}(\boldsymbol{\gamma})$. Then, ${\cal P}_{o}(\boldsymbol{\gamma})$ and ${\cal S}_{o}(\boldsymbol{\gamma},P)$ have the following relations:
@@
Based on Proposition 1, we can solve the MMF problem ${\cal S}_{o}$ using Algorithm 1. In particular, by the relation in (27), ${\cal S}_{o}$ can be solved by solving its inverse QoS problem ${\cal P}_{o}$ iteratively along with a bi-section search for $t^{o}$. However, since the size of the RIS reflection elements $M$ is large, this indirect approach by iteratively solving ${\cal P}_{o}$ can be computationally inefficient.
In the following, we propose a more efficient algorithm to solve ${\cal S}_{o}$ directly. In Section IV-B, PSA is proposed to obtain ${\bf e}$ for the RIS multicast beamforming MMF problem. Below, we propose an alternating PSA (APSA) approach to directly compute a solution for ${\cal S}_{o}$.
V-A Proposed Fast Algorithm
Since problems ${\cal P}_{o}$ and ${\cal S}_{o}$ are the inverse problems, it implies that the optimal structure of the BS multicast beamformer ${\bf w}$ still has the form given in (14). Indeed, using the equivalent channel $\tilde{{\bf h}}_{ik}$ between the BS and the user, the BS mutlicast beamforming is the same as the conventional downlink multicast MMF problem, for which the optimal beamforming structure is [20]
${\bf w}_{i}=\widetilde{{\bf R}}^{-1}(\boldsymbol{\lambda})\widetilde{{\bf H}}_{i}{\bf a}_{i}$ (28)
@@ V-A Proposed Fast Algorithm
where $\widetilde{{\bf R}}={\bf{I}}+\frac{P}{\sigma^{2}}\sum_{i=1}^{G}\sum_{k=1}^{K}\frac{\lambda_{ik}\gamma_{ik}}{\boldsymbol{\lambda}^{T}\boldsymbol{\gamma}}\tilde{{\bf h}}_{ik}\tilde{{\bf h}}_{ik}$.
However, computing $\boldsymbol{\lambda}$ in $\widetilde{{\bf R}}(\boldsymbol{\lambda})$ is not straightforward for the MMF problem. Fortunately, the asymptotic expression of $\widetilde{{\bf R}}(\boldsymbol{\lambda})$ for large $N$ has been obtained in closed-form [20]. We can use it as an approximate expression of $\widetilde{{\bf R}}(\boldsymbol{\lambda})$ to further simplify the computation. Specifically, we rewrite $\tilde{{\bf h}}_{iK}=\sqrt{\tilde{\beta}_{ik}}{\bf g}_{ik}$ where $\tilde{\beta}_{ik}$ represents the large-scale channel variance and $\tilde{{\bf g}}_{ik}\sim{\cal C}{\cal N}(\mathbf{0},{\bf I})$. Then, the asymptotic expression of $\widetilde{{\bf R}}(\boldsymbol{\lambda})$ has the following simple closed form
$\widetilde{{\bf R}}(\boldsymbol{\lambda})\approx{\bf I}+\frac{P\bar{\beta}_{{\bf h}}}{\sigma^{2}K_{\text{tot}}}\sum_{i=1}^{G}\sum_{k=1}^{K_{i}}{\bf g}_{ik}{\bf g}_{ik}^{H}\triangleq\widetilde{{\bf R}}^{\text{\tiny asym}}$ (29)
@@
where $\bar{\beta}_{{\bf h}}\triangleq 1/\Big{(}\frac{1}{K_{\text{tot}}}\sum_{i=1}^{G}\sum_{k=1}^{K_{i}}\frac{1}{\tilde{\beta}_{ik}}\Big{)}$ is the harmonic mean of the large-scale channel variance of all users. Using $\widetilde{{\bf R}}(\boldsymbol{\lambda})$ in (29), we can replace ${\bf w}_{i}$ in (14) with the following expression:
${\bf w}_{i}\approx(\widetilde{{\bf R}}^{\text{\tiny asym}})^{-1}\widetilde{{\bf H}}_{i}{\bf a}_{i}\,,\;i\in{\cal G}.$ (30)
@@
We use ${\bf w}_{i}$ in (30) to develop a low-complexity fast algorithm for ${\cal S}_{o}$. We first convert ${\cal S}_{o}$ into a joint optimization problem w.r.t. $\{{\bf a}_{i}\}$ and ${\bf e}$, given by
@@
where $\widetilde{{\bf C}}_{i}\triangleq(\widetilde{{\bf R}}^{\text{\tiny asym}})^{-1}\widetilde{{\bf H}}_{i}$, ${\bf a}\triangleq[{\bf a}_{1}^{H},\cdots,{\bf a}_{G}^{H}]^{H}$, and the constraints in (26c) are replaced with the equivalent set of constraints in (31c) and (31d) based on (10). Using the similar relaxation technique from ${\cal S}_{{\bf w}}^{\text{eq}}$ to $\widetilde{{\cal S}}_{{\bf w}}$ in (12), we relax $S_{o}^{\text{eq}}$ into the following problem by moving the equality constraint in (31d) into the objective function as a penalty term with a penalty weight $\tilde{\rho}>0$:
$\widetilde{{\cal S}}_{o}:\max_{{\bf a},{\bf e}}$ $\ \bigg{(}\!\min_{i,k}\ \ \frac{\textrm{SINR}_{ik}}{\gamma_{ik}}\bigg{)}+\tilde{\rho}\frac{{\bf e}^{H}{\bf e}}{M}$ (32a)
@@
We note that problem $\widetilde{{\cal S}}_{o}$ is structurally similar to $\widetilde{{\cal S}}_{\bf w}$ in (12), except that $\widetilde{{\cal S}}_{o}$ is a joint optimization problem for $({\bf a},{\bf e})$ while $\widetilde{{\cal S}}_{\bf w}$ is for ${\bf e}$ only. Thus, we propose an APSA approach to alternatingly apply PSA to compute a solution $({\bf a},{\bf e})$ for $\widetilde{{\cal S}}_{o}$.
V-A1 Reformulation
@@
We first convert $\widetilde{{\cal S}}_{o}$ into a min-max problem. Substituting the SINR expression in (1) into the objective function in (32a), for each $k\in{\cal K}_{i}$, $i\in{\cal G}$, we define
@@
$\varphi_{ik}({\bf a},{\bf e})\triangleq$ $-\!\frac{1}{\gamma_{ik}}\frac{|({\bf h}_{ik}^{\text{d}}\!\!+\!{\bf G}_{ik}{\bf e})^{H}\widetilde{{\bf C}}_{i}{\bf a}_{i}|^{2}}{\sum_{j\neq i}|({\bf h}_{ik}^{\text{d}}\!\!+\!{\bf G}_{ik}{\bf e})^{H}\widetilde{{\bf C}}_{j}{\bf a}_{j}|^{2}\!+\!\sigma^{2}}-\tilde{\rho}\frac{{\bf e}^{H}{\bf e}}{M},$ (33)
@@
Following the similar arguments from (15)–(17), let $\boldsymbol{\varphi}({\bf a},{\bf e})$ be a $K_{\text{tot}}\times 1$ vector containing all $\varphi_{ik}({\bf a},{\bf e})$’s, and let ${\bf y}$ denote a $K_{\text{tot}}\times 1$ vector with ${\bf y}\succcurlyeq 0$ and ${\bf 1}^{T}{\bf y}=1$. Then, $\widetilde{{\cal S}}_{o}$ can be transform into the following equivalent problem
$\widetilde{{\cal S}}_{o}^{\prime}:\min_{{\bf a}\in{\cal A},{\bf e}\in{\cal E}}\max_{{\bf y}\in{\cal Y}}$ $\ \tilde{f}({\bf a},{\bf e},{\bf y})$ (34)
@@
where $\tilde{f}({\bf a},{\bf e},{\bf y})\triangleq\boldsymbol{\varphi}^{T}({\bf a},{\bf e}){\bf y}$, ${\cal A}\triangleq\{{\bf a}:\sum_{i=1}^{G}\|\widetilde{{\bf C}}_{i}{\bf a}_{i}\|^{2}\leq P\}$, ${\cal Y}=\{{\bf y}:{\bf y}\succcurlyeq 0,\mathbf{1}^{T}{\bf y}=1\}$, and =${\cal E}=\{{\bf e}:|e_{m}|^{2}\leq 1,\ m\in{\cal M}\}$. All ${\cal A}$, ${\cal E}$, and ${\cal Y}$ are compact convex sets.
V-A2 Alternating Projected Subgradient Algorithm